OpenStudy (anonymous):

Graph the following linear inequality. y<1/9x+9

OpenStudy (anonymous):

@Miracrown

OpenStudy (anonymous):

@charliealbert

OpenStudy (anonymous):

No, it's like this... \[y<\frac{ 1 }{ 9x }+9\]

OpenStudy (anonymous):

I have no idea what that graph means...I just need points or something

OpenStudy (anonymous):

Ohh, sorry I do not know how to graph that, sorry. May be someone else can

Miracrown (miracrown):

If you are in calculus, then we are probably expected to take the derivative to find slopes.

Miracrown (miracrown):

Are you in a calculus class? or precalculus I can give you a general idea of what it looks like.

OpenStudy (anonymous):

No I have algebra or something

Miracrown (miracrown):

The x is in the denominator....correct?

OpenStudy (anonymous):

yes

Miracrown (miracrown):

I assume you're in algebra class, then we need to plot points.

OpenStudy (anonymous):

yes

Miracrown (miracrown):

You are expected to plot points. First we graph y = 1/(9x) + 9 as a dotted line. Basically the graph of y = 1/x looks like this:

Miracrown (miracrown):

 So our final graph will look like this....just a bit different because of the 9's But that is the basic shape.

Miracrown (miracrown):